### v1

# MSegRNN:Enhanced SegRNN Model with Mamba for Long-Term Time Series Forecasting

Abstract
The field of long-term time series forecasting demands handling extensive look-back windows and long-range prediction steps, posing significant challenges for RNN-based methodologies. Among these, SegRNN, a robust RNN-driven model, has gained considerable attention in LTSF analysis for achieving state-of-the-art results while maintaining a remarkably streamlined architecture. Concurrently, the Mamba structure has demonstrated its advantages in small to medium-sized models due to its capability for information selection. This study introduces a variant of SegRNN that preprocesses information using a fine-tuned single-layer Mamba structure. Additionally, it incorporates implicit segmentation and residual structures into the model’s encoding section to further reduce the inherent data iterative cycles of RNN architectures and implicitly integrate inter-channel correlations. This variant, named MSegRNN, utilizes the Mamba structure to select useful information, resulting in a transformed sequence. The linear-strategy-adapted derivative retains the superior memory efficiency of the original SegRNN while demonstrating enhanced performance. Empirical evaluations on real-world LTSF datasets demonstrate the superior performance of our model, thereby contributing to the advancement of LTSF methodologies.
Keywords Long Time Series Analysis  $\cdot$ RNN  $\cdot$ Mamba  $\cdot$ Implicit Segment
1 Introduction
Long-term series forecasting is the process of predicting the future values of a sequence over an extended period. This task primarily aims to forecast long-term trends, which hold significant importance in domains such as climate prediction (Aghelpour et al., 2019), decision support (Nooruldeen et al., 2022), and policy formulation (Zhang et al., 2022). Long-term sequence data often exhibit higher levels of uncertainty, which can result in reduced prediction accuracy. Additionally, longer forecasting horizons require models to incorporate a broader historical context for accurate predictions, thereby increasing the complexity of modeling. Recent advancesments in long-term sequence analysis have shifted towards deep learning methods, highlighting the crucial need to develop diverse and effective time-series methodologies.
Long-term time series analysis and forecasting comprise two critical components: (1) trend identification and analysis, which involves recognizing long-term directional changes in the data, This component is crucial for understanding the behavior of time series, enabling the comprehension of past patterns and the prediction of potential future developments. (2) Uncertainty and noise management: Due to the extended time range of long-term time series data, they often exhibit increased levels of uncertainty and noise. Effectively managing and mitigating these uncertainties and noises is vital for improving the accuracy of predictions.
In the context of long-term time series analysis, it is essential to capture the sequential relationships between data. Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks (Wang et al., 2022; Wu et al., 2022; Foumani et al., 2021), Multilayer Perceptrons (Das et al., 2023; Chen et al., 2023; Zeng et al., 2023), attention-based encoding layers (Nie et al., 2022; Wu et al., 2021; Zhang and Yan, 2022), and Recurrent Neural Networks networks based on sequence structure (Lin et al., 2023). CNN-based methods use convolutional kernels to capture temporal and spatial sequences (Foumani et al., 2021); however, the limited size of these kernels prevent the model from effectively capturing long-range dependencies in time series, thus limiting the expressiveness of the feature representation. Transformer-based methods, leveraging attention mechanisms to improve the capture of spatiotemporal sequences, thereby enhancing feature representation capabilities. MLPs have been exrensively applied in time series prediction, with models such as DLinear (Zeng et al., 2023), which use simple linear layers and channel-independent methods, achieving superior performance compared to the advanced Transformer architectures of its time, prompting a reevaluation of the need for complex Transformer structures in time series prediction.
The Mamba architecture, proposed by Gu and Dao(Gu and Dao, 2023), constitutes a state space model tailored for efficient sequence modeling. It integrates parameterized matrices and a hardware-aware parallel computing algorithm, achieving impressive results in tasks such as language modeling and time series forecasting. Mamba is notable for its ability to capture long-term dependencies while maintaining linear computational complexity, making it a strong competitor to Transformer-based models. This capability is facilitated through its adept management of state space models and the implementation of a selective state space mechanism designed to optimize performance.
RNN-based methods have achieved remarkable success in long-term time series analysis tasks. The inherent sequential architecture of RNNs enables them to handle short-term time series prediction tasks effectively. However, as the length of the predicted sequence increases, the model’s prediction error rapidly escalates with the propagation of features through the RNN recurrent units (Yunpeng et al., 2017). To address this limitation, Shengsheng Lin et al. (Lin et al., 2023) proposed the segmentation-based SegRNN model. This model aims to minimize the number of RNN recurrent iterations while retaining the sequential information, achieving significant success in long-term time series analysis tasks. SegRNN employ a channel-independent strategy to enhance its effectiveness. Nevertheless, this strategy fails to effectively handle noise and capture the correlations between variables. Additionally, even with the adoption of the segmentation strategy, SegRNN still faces the issue of information loss within the recurrent structure.
In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segment strategy. Specifically, we propose a long-term time series prediction model based on Mamba transformation and linear encoding, termed MSegRNN, the main contributions of MSegRNN can be summarized as follows:
  • Introducing the Mamba structure to effectively extract the information
  • Utilizes information between channels to promote the infomation
  • The implementation of residual network in the encoding phase reduces the accumulation of errors during encoding stage.
The structure of this paper is organized as follow: Section 2 formally presents our problem statement. Section 3 discusses related work. Section 4 elaborates on our method of introducing Mamba structure and other technologys. Section 5 describes our experiments and results. Section 6 concludes the paper and suggests future research directions.
2 Problem statement
The analysis of long-term time series involves addressing the challenge of predicting future values of multivariate time series based on historical data. Typically, the problem of time series analysis can be formulated as follows:
$Y=F(X)+\xi$ (1)
In which $X\in\mathbb{R}^{L\times C}$ represents multi-variable time sequences with historical context, $F$ represents the trained neural network structure, $Y\in\mathbb{R}^{H\times C}$ represents the future time sequences to be predicted, $L$ indicates the historical lookback window, $C$ indicates the number of variables, and H indicates the time step to be predicted. Long-term time series analysis aims to expand the potential of predicting time steps, which poses significant challenges for training neural networks.
3 Related Work
A significant amount of work has been dedicated to advancing the development of deep learning in time series analysis. In this section, we introduce literature relevant to our work. These methods can be broadly divided into the following four categories:
3.1 Transformer Models
Transformer models, originally designed for sequence-to-sequence tasks, have achieved significant achievements accross various domains. The core of applying Transformer to long-time series tasks is to use the attention mechanism to capture the long-term dependencies (Hao and Cao, 2020). Extensive research has been devoted to advancing the application of the Transformer architecture in long-time series analysis. Autoformer (Wu et al., 2021) decomposes the sequence into basic blocks within the model, thus gradually decomposing complex time series and enhancing learning ability for complex patterns. Informer (Zhou et al., 2021) constructs a ProbSparse self-attention mechanism, adopting self-attention distillation and a generative decoder to address the high time complexity, large memory consumption, and encoder architecture limitations faced in long-time series analysis. Crossformer (Zhang and Yan, 2022) embeds time series into two-dimensional vector through the Dimension-segment-wise structure, effectively captures dependencies across time and dimensions through the Two-stage-attention structure. The segmentation technique used in Crossformer method is considered beneficial for long time series foresting.
3.2 MLP Models
Extensive research has been dedicated to the application of MLP methods in long-term time series forecasting. Dlinear (Zeng et al., 2023) decomposes time series through the application of moving average kernels and remainder components, yielding two feature variables. These variables are then mapped to the prediction results via linear layers. The Nlinear (Zeng et al., 2023) method adopts a regularization strategy in time series forecasting tasks to enhance performance. Specifically, it subtracts the last value before inputting to the model, mapping through a linear layer, then adding it back for the final prediction. The TiDE (Nie et al., 2022) method introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results. Long-term time series analysis methods based on MLP have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus advancing the development of MLP in long-term time series forecasting.
3.3 CNN Models
The core of time series analysis based on CNN architectures lies in utilizing convolutional kernels to extract temporal and channel features. The Disjoint-CNN (Foumani et al., 2021) approach, while maintaining computational efficiency, enhances prediction accuracy by decomposing one-dimensional convolutional kernels into disjoint temporal and spatial components. Nonetheless, CNN-based methodologies encounter challenges as the forecasting horizon expands. MICN (Wang et al., 2022) aims to fully exploit the latent information in time series by designing a multi-scale hybrid decomposition module to decompose the complex patterns of input time series. Each module within this design leverages subsampling convolution and equidistant convolution to respectively extract local characteristics and global correlations. TimesNet (Wu et al., 2022) transforms one-dimensional time series into a set of two-dimensional tensors based on multiple periodicities. It further introduces the TimesBlock to adaptively uncover these multi-periodic patterns and effectively extracts intricate temporal dynamics from the transformed tensors through parameter-efficient initializations. These methodologies have attained remarkable results in the domain of long-term time series forecasting.
3.4 RNN Models
The RNN architecture was initially designed for processing sequence tasks and excel in short-term time series prediction. However, RNN-based methods face error accumulation issues in long-term time series tasks. SegRNN (Lin et al., 2023) aims to mitigate the accumulation of errors in the recurrent structure while preserving sequence information. Specifically, in the encoding stage, the SegRNN method segments the data before encoding. In the decoding stage, it employs a PMF approach for parallel decoding, significantly enhancing prediction accuracy and the spatiotemporal efficiency.
Nonetheless, the SegRNN method is not without its limitations: (1) SegRNN couldn’t effectivly preporcess the data and extract the inforamtion, which may lead to a rapid accumulation of noise. (2)SegRNN adopts fixed segmentation of the original time series, which may result in partial information loss due to its rigidity. (3) SegRNN aims to minimize the number of recurrent iterations while striving to retain sequential information. However, there are still limitations in minimizing the number of recurrent iterations, which restrict the model’s capability. In the following section, we will propose some strategies to address these issues.
4 Method
To address the challenges faced by current methods, we incorporate mamba structure, implicit segmentation and residual networks into the SegRNN framework. This approach (1) first extracts the original data first to contrain noise, (2)segments the original data through linear mapping to reduce information loss caused by fixed segmentation, and (3) introduces residual structures at the encoding layer to mitigate information loss within the recurrent structure.
We first discuss the overall sturcture that we used in Section 4.1, then we discuss the comparison between explicit segmentation and implicit segmentation in Section 4.2, followed by a discussion on how to introduce residual structures into the model during the encoding process of time series data in Section 4.3.
4.1 Model Structure
The Mamba architecture achieves long-term sequence memory and model selection capabilities by setting the SSM structure parameters as trainable parameters dependent on the sequence content. We have adapted the Mamba architecture to address time series preprocessing problems. The basic structure of our model is shown in Figure 1:
Figure 1: Overall structure of MSegRNN model
Several modifications were applied to the Mamba block and SegRNN model to improve performance and flexibility. Specifically, the convolutional layer in Mamba was removed to enhance computational efficiency, with local information being handled by the SegRNN structure. The SegRNN model’s fixed segmentation was replaced with Implicit Segmentation for adaptability, and residual connections were integrated to strengthen the linkage between segmentation and encoding outputs. Further details on these adjustments are discussed in Section 4.2 and Section 4.3.
4.2 Implicit Segmentation
The explicit segmentation approach employed by SegRNN pre-processes the data such that a given multivariate time series, denoted as $X^{(i)}\in\mathbb{R}^{L\times C}$, where $L$ represents the look-back length and $C$ represents the number of channels in multivariate time series. Upon selecting a segment length $w$, the time series can be transformed into segmented time series $X_{w}^{(i)}\in\mathbb{R}^{n\times w\times C}$, where $n$ represents the number of segments:
$n=\frac{L}{w}$ (2)
Subsequently, the segmented model projects through a linear layer:
$\tilde{X}_{w}^{(i)}=W_{prj}\cdot X_{w}^{(i)}$ (3)
where $W_{prj}\in\mathbb{R}^{w\times d}$ denotes the projection matrix of linear mapping, and $d$ represents the dimensionality of the hidden layer within the RNN structure. Undergoing linear projection transformation, the dimensionality of $\tilde{X}_{w}^{(i)}$ is defined as $\mathbb{R}^{C\times n\times d}$.
However, this segmentation strategy has certain limitations. Once the segment length is chosen, the cut-off points are also determined. This may disrupt the continuity of information in time series analysis, thus limiting the model’s capabilities. The implicit segmentation method aims to break this limitation. Specifically, given a time series analysis vector $X^{(i)}\in\mathbb{R}^{L\times C}$, we first expand its last dimension and then swap the first and second dimensions to obtain $\tilde{X}^{(i)}\in\mathbb{R}^{C\times L\times 1}$ for subsequent processing. Then we utlize a linear layer to transform the last dimesntion of $\tilde{X}^{(i)}$ to get $\bar{X}(i)\in\mathbb{R}^{C\times L\times n}$ to represent the segmented matrix, subsequently, we interchange the second and third dimension, still denoted as $\bar{X}(i)$. Applying linear layer to the last dimension of $\bar{X}(i)$ yields $\bar{X}(i)\in\mathbb{R}^{C\times n\times d}$, which is consistent in dimension with the vector obtained via SegRNN methodology.
Notably, our proposed method does not explicitly define the segmentation length. Instead, it achieves data segmentation and dimensionality transformation through dual linear projections. The incorporation of linear fully connected layers facilitates a more continuous transformation of information. It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe approach but rather represents a general and straightforward technique for information enhancement.
4.3 Residual Structure for SegRNN
SegRNN aims to minimize the number of iterations in the RNN while preserving sequential information. Extensive experiments show that increasing the number of iterations negatively impacts information retention. SegRNN uses data segmentation to reduce the data’s flow through the RNN structure. Specifically, for the vector obtained through segmentation in SegRNN, $\tilde{X}_{w}^{(i)}\in\mathbb{R}^{C\times n\times d}$, we transform it in the RNN structure with a hidden layer dimension $d$ using the following formula:
$\mathbf{h}_{0}$ $=\mathbf{0}$ (4)
$\mathbf{h}_{1}$ $=f(\mathbf{h}_{0},\tilde{\mathbf{X}}_{w}^{(i)}[:,1,:];\mathbf{W})$
$\mathbf{h}_{2}$ $=f(\mathbf{h}_{1},\tilde{\mathbf{X}}_{w}^{(i)}[:,2,:];\mathbf{W})$
$\vdots$
$\mathbf{h}_{n}$ $=f(\mathbf{h}_{n-1},\tilde{\mathbf{X}}_{w}^{(i)}[:,n,:];\mathbf{W})$
where $\mathbf{h}_{t}\in\mathbb{R}^{1\times C\times d}$ represents the hidden state at step $t$, $\tilde{\mathbf{X}}_{w}^{(i)}[:,t,:]$ represents the input vector at step $t$, $\mathbf{W}$ denotes the RNN’s weight parameters, and $f$ is the RNN’s activation function.
Using segmentation techniques, SegRNN reduces the number of iterations from $L$ to $n$. However, the number of iterations still increases with the forecasting time steps. This unavoidably limits the model’s ability due to the loss of information during iterations. Therefore, we introduce a residual structure in the model. For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbb{R}^{C\times L\times n}$, we introduce a linear mapping from $\bar{\mathbf{X}}^{(i)}$ to $\mathbf{h}_{n}$. Specifically, we first merge the last two dimensions and extend the first dimension, obtaining the three-dimensional matrix $\mathbf{X}^{\prime(i)}\in\mathbb{R}^{1\times C\times(L\times n)}$. Then, we transform the last dimension through a linear layer to get $\mathbf{h}^{\prime}_{n}\in\mathbb{R}^{1\times C\times d}$. Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$, obtaining the final output vector from the encoder.
Introducing the residual structure allows some information to bypass the recurrent structure and be directly mapped to the encoder output. This approach minimizes information loss within the recurrent structure, providing a beneficial enhancement for time series data.
5 Experiments
Our experiments are structured as follows. In Section 5.1, we introduce the datasets used for long-term time series analysis. Subsequently, in Section 5.2, we provide a detailed description of the experimental setup and baseline models for comparison. Section 5.3 presents and analyzes the performance metrics. In Section 5.4, we conducted ablation studies to investigate the effectiveness of the model. Finally in Section 5.5, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate the lightweight characteristic of our model. All experiments in this section were conducted on a single NVIDIA 4090 GPU.
5.1 Datasets
The evaluation is conducted across seven real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, and traffic, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings. Detailed dataset information is provided in Table 1.
Table 1: Dataset Basic Attributes
Dataset Electricity Weather ETTh1 ETTh2 ETTm1 ETTm2
Time-Series 321 21 7 7 7 7
Frequency 1 hour 10 mins 1 hour 1 hour 15 mins 15 mins
Time-Points 26,304 52,696 17,420 17,420 69,680 69,680
5.2 Experimental setup and Baselines
The unified configuration of the model is substantially aligned with the SegRNN approach. The look-back window is set to be 96, implicit Segmentation length is set to be 24 or 48 based on the size of datasets. A single layer of Mamba is utilized for preprocessing the data and a single GRU layer is used for sequence processing. The dimensionality of hidden layer with GPU structure is set to be 512, and the training epochs are set to be 30. Dropout rate, learning rate, and batch size vary with the data and the scale of the model.
As baselines, we have selected state-of-the-art and representative models in the long-term time series forecasting domain, comprising the following categories: (i) RNNs: SegRNN; (ii) MLPs: Dlinear; (iii) CNNs: TimesNet; (iv) Transformers: iTransformer, Crossformer, PatchTST.
5.3 Main Result
We selected two evaluation metrics, MSE and MAE, with prediction steps of 96, 192, 336, and 720 time steps. The resulting forecasts are presented in Table 2. Note that all data, except for our method, originate from official code repositories and original papers.
Our method achieved top positions in 37 out of 48 metrics, while the SegRNN method achieved the top position in 7 metrics. The iTransformer, Crossformer, and PatchTST methods each achieved top position in 6, 1 and 1 metrics, respectively. This demonstrates the powerful capabilities of our model.
Table 2: Performance metrics for Long-time series analysis.
Datasets Models RNNs Transformers MLPs CNNs
MSegRNN SegRNN iTransformer Crossformer PatchTST Dlinear TimesNet
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.368 0.383 0.368 0.395 0.386 0.405 0.423 0.448 0.414 0.419 0.386 0.400 0.384 0.402
192 0.422 0.414 0.408 0.419 0.441 0.436 0.471 0.474 0.460 0.445 0.437 0.432 0.436 0.429
336 0.471 0.436 0.444 0.440 0.487 0.458 0.570 0.546 0.501 0.466 0.481 0.459 0.491 0.469
720 0.483 0.459 0.446 0.457 0.503 0.491 0.653 0.621 0.500 0.488 0.519 0.516 0.521 0.500
ETTh2 96 0.275 0.326 0.278 0.335 0.297 0.349 0.745 0.584 0.302 0.348 0.333 0.387 0.340 0.374
192 0.354 0.378 0.359 0.389 0.380 0.400 0.877 0.656 0.388 0.400 0.477 0.476 0.402 0.414
336 0.404 0.416 0.421 0.436 0.428 0.432 1.043 0.731 0.426 0.433 0.594 0.541 0.452 0.452
720 0.407 0.428 0.432 0.455 0.427 0.445 1.104 0.763 0.431 0.446 0.831 0.657 0.462 0.468
ETTm1 96 0.314 0.345 0.330 0.369 0.334 0.368 0.404 0.426 0.329 0.367 0.345 0.372 0.338 0.375
192 0.361 0.374 0.369 0.392 0.377 0.391 0.450 0.451 0.367 0.385 0.380 0.389 0.374 0.387
336 0.388 0.392 0.399 0.412 0.426 0.420 0.532 0.515 0.399 0.410 0.413 0.413 0.410 0.411
720 0.448 0.430 0.454 0.443 0.491 0.459 0.666 0.589 0.454 0.439 0.474 0.453 0.478 0.450
ETTm2 96 0.170 0.248 0.173 0.255 0.180 0.264 0.287 0.366 0.175 0.259 0.193 0.292 0.187 0.267
192 0.233 0.292 0.237 0.298 0.250 0.309 0.414 0.492 0.241 0.302 0.284 0.362 0.249 0.309
336 0.294 0.331 0.296 0.336 0.311 0.348 0.597 0.542 0.305 0.343 0.369 0.427 0.321 0.351
720 0.391 0.389 0.389 0.407 0.412 0.407 1.730 1.042 0.402 0.400 0.554 0.522 0.408 0.403
Electricity 96 0.149 0.242 0.151 0.245 0.148 0.240 0.219 0.314 0.181 0.270 0.197 0.282 0.168 0.272
192 0.165 0.259 0.164 0.258 0.162 0.253 0.231 0.322 0.188 0.274 0.196 0.285 0.184 0.289
336 0.181 0.275 0.180 0.277 0.178 0.269 0.246 0.337 0.204 0.293 0.209 0.301 0.198 0.300
720 0.218 0.307 0.218 0.313 0.225 0.317 0.280 0.363 0.246 0.324 0.245 0.333 0.220 0.320
Weather 96 0.155 0.193 0.165 0.227 0.174 0.214 0.158 0.230 0.177 0.210 0.196 0.255 0.172 0.220
192 0.206 0.244 0.211 0.273 0.221 0.254 0.206 0.277 0.225 0.250 0.237 0.296 0.219 0.261
336 0.260 0.288 0.270 0.318 0.278 0.296 0.272 0.335 0.278 0.290 0.283 0.335 0.280 0.306
720 0.341 0.340 0.357 0.376 0.358 0.349 0.398 0.418 0.354 0.340 0.345 0.381 0.365 0.359
Count 37 7 6 1 1 0 0
5.4 Ablation study
We further conducted an ablation study to verify the effectiveness of incorporating the Mamba structure, implicit segmentation, and residual structures into our model. Using the weather dataset as an example, we performed experiments excluding the Mamba structure and implicit segmentation structure respectivly. For clarity, in the table below, "M" denotes preprocessing using the Mamba structure, and "LR" represents the addition of linear segmentation and residual structure. The results are shown in Table 3:
Table 3: Ablation study on different structure
Datasets Models M&LR&SegRNN LR&SegRNN M&SegRNN SegRNN
MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.155 0.193 0.163 0.202 0.191 0.232 0.165 0.227
192 0.206 0.244 0.207 0.242 0.243 0.277 0.211 0.273
336 0.260 0.288 0.264 0.287 0.305 0.324 0.270 0.318
720 0.341 0.340 0.350 0.342 0.380 0.377 0.357 0.376
Average 0.241 0.266 0.246 0.268 0.280 0.303 0.251 0.298
The ablation study reveals that the model incorporating the Mamba structure, along with implicit segmentation and residual structures performs the best overall. This is followed by the model that includes only implicit segmentation and residual structures. Interestingly, merely adding the Mamba structure does not significantly enhance model performance. This may be due to a synergistic effect between the two structures. Consequently, the ablation study confirms the superiority of our proposed structure.
5.5 Model efficiency
Since our Mamba structure has not undergone hardware-level optimization, this section focuses on comparing the model’s memory usage to verify its lightweight characteristic. The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments. Using the Weather dataset as an example, the memory usage comparison with batch size 64 is shown in Figure 2.
Figure 2: Mean memory consumpation in weather datasets for MSegRNN and SegRNN
As shown in the figure, the model’s exhibits only marginal increases in memory usage compared to SegRNN, affirming its lightweight characteristics.
6 Conclusion
In this work, we present the MSegRNN method: a network that incorporates Mamba structure, implicit segmentation and residual structure into the SegRNN framework. These enhancements mitigate the information loss caused by fixed segmentation in time series while enhance the information from the recurrent structure of the model. It is noteworthy that our proposed implicit segmentation method is not limited to a specific long-term time series forecasting method.
We further conducted experiments on six real-world benchmark datasets for long time series. The results demonstrate that our model yields significant imporvements in some real-world scenarios. To validate the efficiency of our method in terms of space complexity, we compared the memory consumption of MSegRNN against SegRNN on datasets of varying scales. The results suggest that our method maintains high space efficiency, providing a solid foundation for the model’s application.
Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the MSegRNN method excels with corresponding empirical research.