### v4

# ISMRNN: An Implicitly Segmented RNN Method with Mamba for Long-Term Time Series Forecasting

Abstract
Long time series forecasting aims to utilize historical information to forecast future states over extended horizons. Traditional RNN-based series forecasting methods struggle to effectively address long-term dependencies and gradient issues in long time series problems. Recently, SegRNN has emerged as a leading RNN-based model tailored for long-term series forecasting, demonstrating state-of-the-art performance while maintaining a streamlined architecture through innovative segmentation and parallel decoding techniques. Nevertheless, SegRNN has several limitations: its fixed segmentation disrupts data continuity and fails to effectively leverage information across different segments, the segmentation strategy employed by SegRNN does not fundamentally address the issue of information loss within the recurrent structure. To address these issues, we propose the ISMRNN method with three key enhancements: we introduce an implicit segmentation structure to decompose the time series and map it to segmented hidden states, resulting in denser information exchange during the segmentation phase. Additionally, we incorporate residual structures in the encoding layer to mitigate information loss within the recurrent structure. To extract information more effectively, we further integrate the Mamba architecture to enhance time series information extraction. Experiments on several real-world long time series forecasting datasets demonstrate that our model surpasses the performance of current state-of-the-art models.
Keywords Long Time Series Analysis  $\cdot$ RNN  $\cdot$ Mamba  $\cdot$ Implicit Segment
1 Introduction
Long-term series forecasting is the process of predicting the future values of a sequence over an extended period. This task primarily aims to forecast long-term trends, which hold significant importance in domains such as climate prediction (Aghelpour et al., 2019), decision support (Nooruldeen et al., 2022), and policy formulation (Zhang et al., 2022). Long-term sequence data often exhibit higher levels of uncertainty, which can result in reduced prediction accuracy. Additionally, longer forecasting horizons require models to incorporate a broader historical context for accurate predictions, thereby increasing the complexity of modeling. Recent advancesments in long-term sequence analysis have shifted towards deep learning methods, highlighting the crucial need to develop diverse and effective time-series methodologies.
Long-term time series analysis and forecasting comprise two critical components: (1) trend identification and analysis, which involves recognizing long-term directional changes in the data, This component is crucial for understanding the behavior of time series, enabling the comprehension of past patterns and the prediction of potential future developments. (2) Uncertainty and noise management: Due to the extended time range of long-term time series data, they often exhibit increased levels of uncertainty and noise. Effectively managing and mitigating these uncertainties and noises is vital for improving the accuracy of predictions.
In the context of long-term time series analysis, it is essential to capture the sequential relationships between data. Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks (Wang et al., 2022; Wu et al., 2022; Foumani et al., 2021), Multilayer Perceptrons (Das et al., 2023a; Chen et al., 2023; Zeng et al., 2023), attention-based encoding layers (Das et al., 2023b; Nie et al., 2022; Zhang and Yan, 2022), and Recurrent Neural Networks networks based on sequence structure (Lin et al., 2023). CNN-based methods use convolutional kernels to capture temporal and spatial sequences (Foumani et al., 2021); however, the limited size of these kernels prevent the model from effectively capturing long-range dependencies in time series, thus limiting the expressiveness of the feature representation. Transformer-based methods, leveraging attention mechanisms to improve the capture of spatiotemporal sequences, thereby enhancing feature representation capabilities. MLPs have been exrensively applied in time series prediction, with models such as DLinear (Zeng et al., 2023), which use simple linear layers and channel-independent methods, achieving superior performance compared to the advanced Transformer architectures of its time, prompting a reevaluation of the need for complex Transformer structures in time series prediction.
The Mamba architecture, proposed by Gu and Dao(Gu and Dao, 2023), constitutes a state space model tailored for efficient sequence modeling. It integrates parameterized matrices and a hardware-aware parallel computing algorithm, achieving impressive results in tasks such as language modeling and time series forecasting. Mamba is notable for its ability to capture long-term dependencies while maintaining linear computational complexity, making it a strong competitor to Transformer-based models. This capability is facilitated through its adept management of state space models and the implementation of a selective state space mechanism designed to optimize performance.
RNN-based methods have achieved remarkable success in long-term time series analysis tasks. The inherent sequential architecture of RNNs enables them to handle short-term time series prediction tasks effectively. However, as the length of the predicted sequence increases, the model’s prediction error rapidly escalates with the propagation of features through the RNN recurrent units (Yunpeng et al., 2017). To address this limitation, Shengsheng Lin et al. (Lin et al., 2023) proposed the segmentation-based SegRNN model. This model aims to minimize the number of RNN recurrent iterations while retaining the sequential information, achieving significant success in long-term time series foreasting tasks. SegRNN employs a segmented approach to time series, which may disrupt the continuity of the sequences. Although the segmentation strategy effectively balances sequence information and iteration count, information loss within the recurrent structure occurs as the prediction horizon extends, thereby limiting the model’s capability.
In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segmentation strategy. Specifically, we propose a long-term time series forecasting model that utilizes the Mamba structure, implicit segmentation, and residual blocks, termed ISMRNN, the main contributions of ISMRNN can be summarized as follows:
  • Introducing the Mamba structure to effectively preprocessing the information.
  • Utilizing implicit segmentation for denser information exchange during the segmentation phase.
  • The implementation of residual network in the encoding phase reduces the accumulation of errors during encoding stage.
2 Problem Statement
The analysis of long-term time series involves addressing the challenge of predicting future values of multivariate time series based on historical data. Typically, the problem of time series analysis can be formulated as follows:
$Y=F(X)+\xi$ (1)
In which $X\in\mathbf{R}^{L\times C}$ represents multivariable time sequences with historical context, $F$ represents the trained neural network structure, $Y\in\mathbf{R}^{H\times C}$ represents the future time sequences to be predicted, $L$ indicates the historical lookback window, $C$ indicates the number of variables, and $H$ indicates the time step to be predicted. Long-term time series analysis aims to expand the potential of predicting time steps, which poses significant challenges for training neural networks.
A significant amount of work has been dedicated to advancing the development of deep learning in time series analysis. These methods can be broadly divided into the following categories:
2.1 Transformer Models
The core of applying Transformer models to long-term time series tasks lies in their ability to capture long-term dependencies using the attention mechanism (Hao and Cao, 2020). The general form of the self-attention mechanism is given by:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (2)
Significant research efforts have advanced the application of Transformer architectures in long-term time series analysis. For instance, PatchTST(Nie et al., 2022) introduces a method of segmenting time series into patches, reducing computational complexity and improving the model’s ability to capture local semantic information. It also employs self-supervised learning, masking random patches and training the model to reconstruct them, enhancing forecasting accuracy. iTransformer(Liu et al., 2023) addresses scalability by integrating hierarchical attention mechanisms and temporal convolutional networks, capturing both short-term and long-term dependencies more effectively. This lightweight architecture balances accuracy and efficiency, making it suitable for real-world applications. Crossformer (Zhang and Yan, 2022) employs a Dimension-segment-wise structure to embed time series into a two-dimensional vector, effectively capturing dependencies across time and dimensions through its Two-stage-attention structure. The segmentation technique utilized in Crossformer is particularly advantageous for long-term time series forecasting.
2.2 MLP Models
MLP methods have garnered significant attention in long-term time series forecasting due to their robust performance. Dlinear (Zeng et al., 2023) decomposes time series through the application of moving average kernels and remainder components, yielding two feature variables. These variables are then mapped to the prediction results via linear layers. The Dlinear model can be described as:
$\hat{Y}=W_{1}(MA(X))+W_{2}(R(X))$ (3)
where $MA(X)$ represents the trend component obtained by the moving average kernel, $R(X)$ represents the remainder component, and $W_{1}$, $W_{2}$ are the weights of the linear layers.
TiDE (Das et al., 2023b) introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results. The core operation of TiDE involves projecting features and encoding them through residual blocks, which can be described as:
$\hat{Y}=\text{ResBlock}(\text{ReLU}(W_{f}X+b_{f})+\text{Covariates})$ (4)
where $W_{f}$ and $b_{f}$ are the weights and biases of the projection layer, respectively. The ResBlock operation encapsulates the residual connections used to maintain the integrity of the encoded features while integrating the covariates.
MLP-based methods for long-term time series analysis have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus significantly advancing the development of MLP in long-term time series forecasting.
2.3 CNN Models
The core of time series analysis based on CNN architectures lies in utilizing convolutional kernels to extract temporal and channel features. MICN (Wang et al., 2022) aims to fully exploit the latent information in time series by designing a multi-scale hybrid decomposition module to decompose the complex patterns of input time series. Each module within this design leverages subsampling convolution and equidistant convolution to respectively extract local characteristics and global correlations. TimesNet (Wu et al., 2022) transforms one-dimensional time series into a set of two-dimensional tensors based on multiple periodicities. It further introduces the TimesBlock to adaptively uncover these multi-periodic patterns and effectively extracts intricate temporal dynamics from the transformed tensors through parameter-efficient initializations. These methodologies have attained remarkable results in the domain of long-term time series forecasting.
2.4 SegRNN Model
As the only RNN-based model specifically designed for long-term time series tasks, SegRNN has achieved remarkable performance. SegRNN addresses the challenge of retaining sequence information while minimizing information loss within the recurrent structure through two key methods:
  • SegRNN replaces point-wise iterations with segment-wise iterations. This approach balances the trade-off between sequence information retention and the number of recurrent iterations, leading to improved outcomes.
  • SegRNN replaces recurrent multi-step forecasting with parallel multi-step forecasting. This parallel decoding approach enhances the model’s predictive capability and inference speed.
By combining these two strategies, SegRNN achieves excellent performance within a lightweight model structure, along with superior inference speed.
3 Method
3.1 Overview
Previous work, such as SegRNN, has achieved remarkable results by employing segmentation and parallel decoding techniques. However, the fixed segmentation approach restricts the model’s capacity and disrupts data continuity. Additionally, segmentation methods still face the challenge of cumulative information processing within the recurrent structure.
In this paper, we propose the ISMRNN method to overcome these limitations through three strategies, as illustrated in Figure 1. We first introduce implicit segmentation, which allows for dense information exchange during the segmentation stage. Subsequently, we incorporate a residual structure to enable partial information bypassing the recurrent structure, thereby reducing information loss in recurrent structure. We also integrate the Mamba architecture to selectively process time series information, further enhancing the model’s capability in handling temporal data.
Figure 1: Overall structure of ISMRNN model
3.2 Mamba structure
The Mamba architecture, initially proposed by Gu and Dao (2023), has demonstrated outstanding performance in various fields such as large language models and long time series forecasting. By setting model parameters to be content-related, the Mamba architecture achieves powerful data selection capabilities, which are crucial for time series forecasting tasks. Therefore, we utilize the Mamba architecture at the initial stage to preprocess and enhance our time series data. It is important to note that our model does not incorporate the convolutional layer of the Mamba architecture except for weather dataset. This difference is mainly based on performance considerations, as we believe that local information can be effectively captured through segmentation.
3.3 Implicit Segmentation
The segmentation method employed by SegRNN, which we refer to as explicit segmentation to distinguish it from our segmentation approach, is implemented by truncating the time dimension. For a given time series $X^{(i)}\in\mathbf{R}^{L\times C}$, selecting a segmentation length $w$ results in segmented vectors of size $X_{w}^{(i)}\in\mathbf{R}^{n\times w\times C}$, where $n$ represents the number of segments.
$n=\frac{L}{w}$ (5)
Subsequently, we use linear mapping to transform the dimension $w$ into $d$, obtaining the embedded hidden state of the segmented vectors $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$. This segmentation method retains the sequence information while reducing the iteration count within the recurrent structure. However, this approach has some drawbacks. The simple truncation of data disrupts continuity, and there is no effective exchange of information between different segments, thus limiting the model’s capability.
To address these issues, we propose an implicit segmentation strategy, as illustrated in Figure 2. This strategy initially decomposes the time series through linear mapping and subsequently transforms the information into the embedded hidden state $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$. Notably, after segmentation, each segment includes both the information of that segment and some additional information, ensuring more continuous information processing and thereby enhancing the model’s performance.
Figure 2: Comparsion of SegRNN Segmentation with Implicit Segmentation, we utilize a linear layers, from 1 to $n$, to decompose the time series. Subsequently, a second mapping is applied to compress the dimensions. Note that each segmented vector embedding can encompass information from other segments of the original time series.
Furthermore, our proposed method does not explicitly define the segmentation length. Instead, it achieves data segmentation and dimensionality transformation through dual linear projections. The incorporation of linear fully connected layers facilitates a more continuous and dense transformation of information. It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe model but rather represents a general and straightforward technique for information enhancement.
3.4 Residual Structure
SegRNN aims to minimize the number of iterations in the RNN while preserving sequential information. Extensive experiments show that increasing the number of iterations negatively impacts information retention. SegRNN uses data segmentation to reduce the data’s flow through the RNN structure. Specifically, for the vector obtained through segmentation in SegRNN, $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$, we transform it in the RNN structure with a hidden layer dimension $d$ using the following formula:
$\mathbf{h}_{0}$ $=\mathbf{0}$ (6)
$\mathbf{h}_{1}$ $=f(\mathbf{h}_{0},\tilde{\mathbf{X}}_{w}^{(i)}[:,1,:];\mathbf{W})$
$\mathbf{h}_{2}$ $=f(\mathbf{h}_{1},\tilde{\mathbf{X}}_{w}^{(i)}[:,2,:];\mathbf{W})$
$\vdots$
$\mathbf{h}_{n}$ $=f(\mathbf{h}_{n-1},\tilde{\mathbf{X}}_{w}^{(i)}[:,n,:];\mathbf{W})$
where $\mathbf{h}_{t}\in\mathbf{R}^{1\times C\times d}$ represents the hidden state at step $t$, $\tilde{\mathbf{X}}_{w}^{(i)}[:,t,:]$ represents the input vector at step $t$, $\mathbf{W}$ denotes the RNN’s weight parameters, and $f$ is the RNN’s activation function.
Using segmentation techniques, SegRNN reduces the number of iterations from $L$ to $n$. However, the number of iterations still increases with the forecasting time steps, which limits the model’s capability due to the loss of information during these iterations. To address this issue, we incorporate a residual structure into the model. For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbf{R}^{C\times L\times n}$, we apply a linear mapping to $\bar{\mathbf{X}}^{(i)}$, resulting in $\mathbf{h}^{\prime}_{n}$. Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$ to obtain the final output vector from the encoder.
Introducing the residual structure allows some information to skip the recurrent structure and be directly mapped to the encoder output. This approach minimizes information loss within the recurrent structure, providing a beneficial enhancement for time series data.
4 Experiments
Our experiments are structured as follows. We first introduce the datasets used for long-term time series analysis. Subsequently, we provide a detailed description of the experimental setup and baseline models for comparison. Then we presents and analyzes the performance metrics. We further conducted ablation studies to investigate the effectiveness of the model. Finally, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate its lightweight characteristic. All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/ISMRNN.git.
4.1 Datasets
The evaluation is conducted across seven real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings. Detailed dataset information is provided in Table 1.
Table 1: Dataset Basic Attributes
Data Electricity Weather ETTh1 ETTh2 ETTm1 ETTm2
Chan. 321 21 7 7 7 7
Freq. 1 hour 10 mins 1 hour 1 hour 15 mins 15 mins
Points 26,304 52,696 17,420 17,420 69,680 69,680
4.2 Experimental setup and Baselines
The unified configuration of the model is substantially aligned with the SegRNN approach. The look-back window is set to be 96, a single layer of Mamba is utilized for preprocessing the data and a single GRU layer is used for sequence processing. The dimensionality of hidden layer with GPU structure is set to be 512, and the training epochs are set to be 30. Dropout rate, learning rate, and batch size vary with the data and the scale of the model.
As baselines, we have selected state-of-the-art and representative models in the long-term time series forecasting domain, comprising the following categories: (i) RNNs: SegRNN; (ii) MLPs: Dlinear; (iii) CNNs: TimesNet; (iv) Transformers: iTransformer, Crossformer, PatchTST.
4.3 Main Result
We selected two evaluation metrics, MSE and MAE, with prediction steps of 96, 192, 336, and 720 time steps. The resulting forecasts are presented in Table 2. Note that all data, except for our method, originate from official code repositories and original papers.
Our method achieved top positions in 36 out of 48 metrics, while the SegRNN method and the iTransformer, each achieved top position in 7 and 6 metrics, respectively. This demonstrates the powerful capabilities of our model.
Table 2: Performance metrics for Long-time series analysis.
Datasets RNNs Transformers MLPs CNNs
ISMRNN SegRNN iTransformer Crossformer PatchTST Dlinear TimesNet
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.365 0.384 0.368 0.395 0.386 0.405 0.423 0.448 0.414 0.419 0.386 0.400 0.384 0.402
ETTh1 192 0.415 0.413 0.408 0.419 0.441 0.436 0.471 0.474 0.460 0.445 0.437 0.432 0.436 0.429
336 0.463 0.441 0.444 0.440 0.487 0.458 0.570 0.546 0.501 0.466 0.481 0.459 0.491 0.469
720 0.468 0.460 0.446 0.457 0.503 0.491 0.653 0.621 0.500 0.488 0.519 0.516 0.521 0.500
96 0.275 0.326 0.278 0.335 0.297 0.349 0.745 0.584 0.302 0.348 0.333 0.387 0.340 0.374
ETTh2 192 0.354 0.378 0.359 0.389 0.380 0.400 0.877 0.656 0.388 0.400 0.477 0.476 0.402 0.414
336 0.404 0.416 0.421 0.436 0.428 0.432 1.043 0.731 0.426 0.433 0.594 0.541 0.452 0.452
720 0.407 0.428 0.432 0.455 0.427 0.445 1.104 0.763 0.431 0.446 0.831 0.657 0.462 0.468
96 0.314 0.345 0.330 0.369 0.334 0.368 0.404 0.426 0.329 0.367 0.345 0.372 0.338 0.375
ETTm1 192 0.361 0.374 0.369 0.392 0.377 0.391 0.450 0.451 0.367 0.385 0.380 0.389 0.374 0.387
336 0.388 0.392 0.399 0.412 0.426 0.420 0.532 0.515 0.399 0.410 0.413 0.413 0.410 0.411
720 0.448 0.430 0.454 0.443 0.491 0.459 0.666 0.589 0.454 0.439 0.474 0.453 0.478 0.450
96 0.170 0.248 0.173 0.255 0.180 0.264 0.287 0.366 0.175 0.259 0.193 0.292 0.187 0.267
ETTm2 192 0.233 0.292 0.237 0.298 0.250 0.309 0.414 0.492 0.241 0.302 0.284 0.362 0.249 0.309
336 0.294 0.331 0.296 0.336 0.311 0.348 0.597 0.542 0.305 0.343 0.369 0.427 0.321 0.351
720 0.391 0.389 0.389 0.407 0.412 0.407 1.024 0.679 0.394 0.390 0.387 0.370 0.400 0.388
96 0.149 0.242 0.151 0.245 0.148 0.240 0.219 0.314 0.181 0.270 0.197 0.282 0.168 0.272
Electricity 192 0.165 0.259 0.164 0.258 0.162 0.253 0.231 0.322 0.188 0.274 0.196 0.285 0.184 0.289
336 0.181 0.275 0.180 0.277 0.178 0.269 0.246 0.337 0.204 0.293 0.209 0.301 0.198 0.300
720 0.218 0.307 0.218 0.313 0.225 0.317 0.280 0.363 0.246 0.324 0.245 0.333 0.220 0.320
96 0.147 0.187 0.165 0.227 0.174 0.214 0.158 0.230 0.177 0.210 0.196 0.255 0.172 0.220
Weather 192 0.198 0.238 0.211 0.273 0.221 0.254 0.206 0.277 0.225 0.250 0.237 0.296 0.219 0.261
336 0.256 0.282 0.270 0.318 0.278 0.296 0.272 0.335 0.278 0.290 0.283 0.335 0.280 0.306
720 0.334 0.337 0.357 0.376 0.358 0.349 0.398 0.418 0.354 0.340 0.345 0.381 0.365 0.359
Count 36 7 6 0 0 0 0
We subsequently conducted experiments to observe how the model’s capability varies with changes in the look-back window. The hyperparameter settings of the method were kept consistent with look-back window 96, and the horizon was uniformly set to 336. The evaluation was performed on the ETTh2 dataset, with the MSE and MAE losses of the method shown in the Figure 3:
Figure 3: Comparsion of SegRNN and ISMRNN with vary look-back window length
Our model exhibited smaller losses across various look-back windows, demonstrating the superiority of our method. Notably, the advantage of our method is more pronounced with smaller look-back windows. As the look-back window increases, this advantage diminishes, possibly due to noise introduced by segmentation affecting the model. Nevertheless, our model maintained a considerable advantage.
4.4 Ablation study
We further conducted an ablation study to verify the effectiveness of incorporating the Mamba structure, implicit segmentation, and residual structures into our model. We performed experiments excluding the Mamba structure and implicit segmentation structure respectivly. For clarity, in the table below, "M" denotes preprocessing using the Mamba structure, and "LR" represents the addition of linear segmentation and residual structure. The results are shown in Table 3:
Table 3: Ablation study on different structure
Datasets Models M&LR&SegRNN LR&SegRNN M&SegRNN SegRNN
MSE MAE MSE MAE MSE MAE MSE MAE
96 0.147 0.187 0.163 0.202 0.191 0.232 0.165 0.227
Weather 192 0.199 0.238 0.207 0.242 0.243 0.277 0.211 0.273
336 0.256 0.282 0.264 0.287 0.305 0.324 0.270 0.318
720 0.332 0.337 0.350 0.342 0.380 0.377 0.357 0.376
96 0.365 0.384 0.377 0.387 0.480 0.446 0.368 0.395
ETTh1 192 0.415 0.413 0.423 0.416 0.520 0.472 0.408 0.419
336 0.463 0.441 0.475 0.437 0.543 0.491 0.444 0.440
720 0.468 0.460 0.476 0.454 0.549 0.515 0.446 0.457
96 0.275 0.326 0.281 0.327 0.326 0.368 0.278 0.335
ETTh2 192 0.354 0.378 0.368 0.385 0.418 0.420 0.359 0.389
336 0.404 0.416 0.402 0.416 0.462 0.456 0.421 0.436
720 0.407 0.428 0.421 0.435 0.467 0.472 0.432 0.455
96 0.314 0.345 0.315 0.346 0.535 0.454 0.330 0.369
ETTm1 192 0.361 0.374 0.362 0.376 0.556 0.469 0.369 0.392
336 0.388 0.392 0.390 0.395 0.576 0.480 0.399 0.412
720 0.448 0.430 0.446 0.427 0.631 0.509 0.454 0.443
96 0.170 0.248 0.173 0.250 0.215 0.298 0.173 0.255
ETTm2 192 0.233 0.292 0.235 0.293 0.283 0.339 0.237 0.298
336 0.294 0.331 0.297 0.332 0.352 0.381 0.296 0.336
720 0.391 0.389 0.392 0.388 0.463 0.441 0.389 0.407
Average 0.334 0.355 0.341 0.357 0.425 0.411 0.340 0.372
4.5 Model efficiency
This section is dedicated to evaluating the model’s time and memory efficiency to substantiate its lightweight nature. The experiments are conducted on weather datasets because the other datasets utilize the Mamba architecture are implemented without convolutional layers and hardware optimization. The lightweight characteristic is primarily due to the use of a small number of parameters in our Mamba structure experiments. Using the Weather dataset as an example, the time and memory usage comparison with batch size 8 conducted on a single NVIDIA T4 GPU is shown in Figure 4.
Figure 4: Comparsion of SegRNN and ISMRNN with time and space efficiency.
As shown in the figure, compared to SegRNN, our model only increases memory usage and training time. However, in contrast to some transformer-based models, our method still holds significant spatiotemporal advantages.
5 Conclusion
In this work, we present the ISMRNN method: a network that incorporates Mamba structure, implicit segmentation and residual structure into the SegRNN framework. These enhancements mitigate the information loss caused by fixed segmentation in time series while enhance the information from the recurrent structure of the model. It is noteworthy that our proposed implicit segmentation method is not limited to a specific long-term time series forecasting method.
We further conducted experiments on six real-world benchmark datasets for long time series. The results demonstrate that our model yields significant imporvements in some real-world scenarios. To validate the efficiency of our method in terms of space complexity, we compared the memory consumption of ISMRNN against SegRNN on datasets of varying scales. The results suggest that our method maintains high space efficiency, providing a solid foundation for the model’s application.
Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the ISMRNN method excels with corresponding empirical research.
Appendix
Appendix A Datasets
For our Long-term Time Series Forecasting (LTSF) study, we employed several widely-used multivariate datasets, including:
ETTs: This collection consists of electricity transformer data from two counties in China, spanning two years. The datasets include various subsets: ETTh1 and ETTh2 provide hourly data, and ETTm1 and ETTm2 offer 15-minute interval data. Each entry encompasses the target variable "oil temperature" alongside six power load indicators.
Electricity: This dataset contains hourly electricity consumption records for 321 clients from 2012 to 2014, offering insights into consumer energy usage patterns over time.
Weather: This dataset captures 21 meteorological variables, such as temperature and humidity, at 10-minute intervals throughout 2020. It serves as a comprehensive resource for weather prediction and climate studies.
These datasets are instrumental for benchmarking and advancing forecasting models in the context of long-term time series analysis.
Appendix B Code Framework
The implementation of our model framework is primarily derived from the official SegRNN code and the implementation of Mamba. This includes data preprocessing and evaluation, model parameter settings and device selection, as well as result output. We sincerely thank the authors of these open-source frameworks and methods for their contributions to our experiments.
The official SegRNN source code is available at: https://github.com/lss-1138/SegRNN.
The official minimal implementation of Mamba can be found at:https://github.com/johnma2006/mamba-minimal.
The official implementation of Mamba can be found at:https://github.com/state-spaces/mamba.git
Appendix C Detail Configuration
The main hyperparameter settings of our model are shown in Table 4. The parameter $d_{\text{state}}$ is a hyperparameter in the Mamba architecture, which is detailed in Mamba [Gu and Dao, 2023]. The parameter settings and results for other methods are taken from their original papers or official code repositories.
Table 4: Detail configuration of ISMRNN
Datasets d_state seg_len use_conv learning_rate dropout
ETTh1 2 12 False 0.0003 0.1
ETTh2 4 24 False 0.0015 0.5
ETTm1 4 24 False 0.0010 0.1
ETTm2 4 24 False 0.0010 0.1
Weather 4 24 True 0.0007 0.1
Electricity 2 48 False 0.0012 0.01
We also employed a different learning rate adjustment strategy. The learning rate remained constant for the first 15 epochs and then gradually multiplied by 0.9. This approach actually increases the learning rate compared to the adjustment method used in SegRNN.
Appendix D Comparison with extra methods
We observed that some methods also utilized Mamba for information augmentation, we selected several long-term time series analysis models based on the Mamba architecture: Time-Machine[Ahamed and Cheng, 2024], Bi-Mamba4TS[Liang et al., 2024], and DT-Mamba[Wu et al., 2024], for comparison. The results are from official paper, respectivly . The result is shown in Table 5.
Table 5: Performance of different models
Datasets Models ISMRNN TimeMachine Bi-Mamba4TS DTMamba
MSE MAE MSE MAE MSE MAE MSE MAE
96 0.147 0.187 0.164 0.208 0.164 0.212 0.171 0.218
Weather 192 0.198 0.238 0.211 0.250 0.214 0.256 0.220 0.257
336 0.256 0.282 0.256 0.290 0.269 0.296 0.274 0.296
720 0.334 0.337 0.342 0.343 0.348 0.349 0.349 0.346
96 0.365 0.384 0.364 0.387 0.376 0.403 0.386 0.399
ETTh1 192 0.415 0.414 0.415 0.416 0.411 0.425 0.426 0.424
336 0.463 0.441 0.429 0.421 0.455 0.445 0.480 0.450
720 0.468 0.460 0.458 0.453 0.460 0.464 0.484 0.470
96 0.275 0.326 0.275 0.334 0.289 0.341 0.290 0.340
ETTh2 192 0.354 0.378 0.349 0.381 0.367 0.389 0.366 0.392
336 0.404 0.416 0.340 0.381 0.410 0.424 0.380 0.409
720 0.407 0.428 0.411 0.433 0.421 0.439 0.416 0.437
96 0.314 0.345 0.317 0.355 0.312 0.354 0.325 0.360
ETTm1 192 0.361 0.374 0.357 0.378 0.358 0.383 0.375 0.386
336 0.388 0.392 0.379 0.399 0.388 0.404 0.396 0.405
720 0.448 0.430 0.445 0.436 0.444 0.433 0.454 0.442
96 0.170 0.248 0.175 0.256 0.174 0.259 0.177 0.259
ETTm2 192 0.233 0.292 0.239 0.299 0.240 0.304 0.240 0.300
336 0.294 0.331 0.287 0.332 0.303 0.345 0.310 0.345
720 0.391 0.389 0.371 0.385 0.402 0.407 0.395 0.394
96 0.149 0.242 0.142 0.236 0.142 0.238 0.166 0.256
Electricity 192 0.165 0.259 0.158 0.250 0.157 0.253 0.178 0.268
336 0.181 0.275 0.172 0.268 0.172 0.271 0.197 0.289
720 0.218 0.307 0.207 0.298 0.200 0.297 0.243 0.326
It can be observed that our method performs notably well on the weather, ETTh2, and ETTm2 datasets, while the TimeMachine method shows significant advantages on the ETTh1 and Electricity datasets. Additionally, the Bi-Mamba4TS method also demonstrates a clear advantage on the Electricity dataset.
Appendix E Prediction Result
We subsequently observed the partial prediction results of the model using the weather dataset, with a look-back window of 96 and a prediction horizon of 336. The results are shown in Figure 5:
Figure 5: Prediction result of different method. The four methods are Dlinear, ISMRNN, SegRNN, and PatchTST in sequence. Intuitively, our method also has the best fitting performance, which is consistent with the fact that ISMRNN has the minimum loss among the models.
Appendix F Ablation study on using convolution layers
In the main text’s model structure, we removed the convolutional layer from the Mamba structure due to performance issues. We hypothesized that local features might have been redundantly captured. To investigate the impact of the convolutional layer on the model performence, we conducted experiments. The results are shown in Table 6.
Table 6: Ablation study on convlution layer
Datasets Models WithoutConv WithConv
MSE MAE MSE MAE
96 0.365 0.384 0.412 0.426
ETTh1 192 0.415 0.414 0.444 0.443
336 0.463 0.441 0.477 0.460
720 0.468 0.460 0.502 0.492
96 0.275 0.326 0.282 0.338
ETTh2 192 0.354 0.378 0.372 0.396
336 0.404 0.416 0.406 0.424
720 0.407 0.428 0.430 0.456
96 0.314 0.345 0.323 0.355
ETTm1 192 0.361 0.374 0.373 0.384
336 0.388 0.392 0.397 0.401
720 0.448 0.430 0.452 0.432
96 0.170 0.248 0.180 0.264
Ettm2 192 0.233 0.292 0.240 0.305
336 0.294 0.331 0.299 0.343
720 0.391 0.389 0.396 0.399
It can be observed that except for the Weather dataset in our experiments, the convolutional layer may interfere with the model’s predictions in most scenarios, thereby limiting the model’s capability.